The driver of a car, weighing 10,000 N is travelling 30 m/s when she sees an N is obstacle on the road and brakes. The coefficient of friction between her wheels and the road is 0.8. She stops in 10 s. How much work, in J, is done by the friction? Select one: a. 1,200,000 v b. cant tell C. 14,000 d. 120,000 e. 1,400,000 The correct answer is: 1,200,000 To determine how well a bicycle wheel is lubricated, the mechanic in the repair shop starts it spinning and lets it coast, measuring the time until it stops. She counts the number of revolutions to be 20 in 60 s. What is the angular deceleration?
Answer #1

\alpha = \omega / t

\omega = 2\pi \nu = 2\pi/T

So, angular speed =2\pi 20/ 60 = 2\pi/3

Deceleration = 2\pi/(3 x 60) = pi/ 90 rad /s^2